# A Survey on Automatic Credibility Assessment of Textual Credibility Signals in the Era of Large Language Models

### v1

@@ 9.2. Dataset availability and multilinguality
Dataset availability heavily differs across various categories of credibility signals. Firstly, we observed a significant lack of large-enough datasets providing overall (expertly-determined) annotation of credibility that can be used for training and evaluating solutions on credibility assessment task (Section 4.1). Furthermore, datasets containing annotations for overall credibility and (at least some) credibility signals at the same time are even more scarce. As a result, many researchers opted to use (more available) fake news datasets as a replacement. At this place, we would like to highlight again that fake news annotations cannot reliably replace credibility annotations (non-credible content is not necessarily only false content, since credibility is a broader concept going beyond content truthfulness).
@@
On the other hand, the situation with dataset availability for individual categories of credibility signals is much better. This is especially thanks to data challenges (particularly SemEval tasks and CLEF CheckThat! Labs) in which either new datasets were introduced or existing datasets were extended (with more data, new languages, or additional types of annotations). Unfortunately, in some cases (e.g., (Piskorski et al., 2023)), the datasets from data challenges are not published completely and a hidden/testing set is not shared with researchers not even when the competition is over (and thus only the training/validation sets remain available for further research).
@@
Especially for international news or the investigation of global claims, journalists and fact-checkers need to verify the credibility of information by cross-checking sources in different languages. Furthermore, emerging disinformation in one country could be spread to other countries, especially when its topic is global (e.g., pandemic, wars, international relations). Therefore, it is important to implement credibility analysis tools that support multiple languages. However, due to the scarcity of multilingual datasets, trained models can exhibit biases towards some languages and cultures and hence can underperform on content in low-resourced languages.
@@
To overcome this issue, global collaborations could be initiated for creating multilingual resources. In this direction, we can already observe a positive trend within data challenges. Many of them introduced multilingual datasets, commonly considering some languages as surprise ones (i.e., languages that are present in a test set, but missing in a train/validation set). Such approach motivates participants to develop multilingual solutions that are capable to make a prediction in a zero-shot setting (considering a language a predicted content is written in).
@@
Besides datasets availability and multilinguality, a quality of annotations remain another challenge. Annotation of overall credibility as well as individual credibility signals is many times highly subjective (as also shown in (Bhuiyan et al., 2020)), especially in cases such as persuasion techniques where presence of signal and borders between various signals can be blurred. The situation is getting even more challenging in multilingual settings, where typically native speakers are needed to annotate data. Firstly, acquiring human experts fluent in several languages is challenging itself. Secondly, organizing and consolidating annotation process (including post-annotation verification) is a complex task. Considering also our own experience working with such datasets, the provided labels cannot be easily verified and many times we identified (inevitable) incorrect labels.
@@
Last but not least, credibility assessment and automatic detection of credibility signals naturally happen in very dynamically evolving (online) environment. New topics and global events constantly emerge, causing significant data and concept drifts. In some cases, such drifts can cause that the existing datasets may become obsolete and non-representative. Secondly, the list of credibility signals evolve itself. We can take a machine-generated text as an illustrative example. This kind of signal become highly relevant only recently with generative LLMs becoming easily available for a large end user base (and unfortunately also bad actors). Such new/redefined signals thus naturally result into a demand for new datasets. Finally, dynamics of this area also lies in its adversarial character. Bad actors (e.g., the ones who are spreading propaganda or disinformation) will always try to get their content undetected as low-credible one by employing various obfuscation techniques. This must be taken into consideration when introducing new datasets. By continuing with an illustrative example of machine-generated text, there are already datasets (e.g., (Macko et al., 2024)) providing besides machine-generated texts also their alternative versions after applying several authorship obfuscation techniques, what allows to train and evaluate more robust detection models.
@@ 9.3. Ethical and legal issues
Credibility assessment of online (primarily social media) content is from its nature an area that must address several ethical and legal issues. Firstly, similarly to other related research areas (e.g., on fake news detection), it may be potentially misused by bad actors in order to create a content that appears to be more credible. Also by following open-research spirit, publishing credibility assessment systems can theoretically result into misusing such system in the adversarial manner to tune disinformation/propaganda generation systems and allowing them to stay undetected. This kind of potential threat is, however, an analogical issue to the security domain and the principle of security by obscurity. Nothing prevents bad actors to develop their own credibility assessments systems and apply then in adversarial training scenario. Moreover, positive outcomes of credibility assessment research are more tangible, with many practical tools (as also showed across this survey) already put into the hands of media professionals or general public.
@@
Besides potential misuse, additional ethical considerations must be addressed thoroughly during the research activities. First of all, training various classification/detection systems is inherently a subject of potential biases. Such biases can come directly from the datasets used for the training purposes (in terms of data selection, data pre-/post-processing, or data annotation itself) or can be introduced during training the models, especially when fine-tuning pre-trained LLMs that have incorporated biases by themselves (including biases between high- and low-resource languages). Secondly, the authors should always clearly formulate intended use and failure cases of the trained models/deployed tools. In this way, we may prevent media professionals to over-rely on the predicted (potentially incorrect) values. Fortunately, the two-step approach to credibility assessment (i.e., detect more granular credibility signals and then aggregate them) makes the whole process more transparent. Last but not least, explainability and interpretability of the models’ predictions plays a crucial role in this area, since credibility assessment must be credible itself, otherwise it would not provide expected level of trust to its end users.
@@
Besides ethical issues, the research in this area must address also several legal issues, many of them are shared with other research works on social media. Firstly, as we also witness recently, social media platforms can change their data processing policies and restrict access to their data, which may delay progress in research and development of credibility tools. Additionally, social media data must be anonymized to protect user privacy before being used as training data or for the model inference. Media organizations also impose limitations due to copyright laws, with some not permitting their content to be used in AI tool development. LLMs, especially closed LLMs such as ChatGPT and GPT-4, lack transparency regarding pre-training data and LLMs can memorize content in their pretraining data (Karamolegkou et al., 2023; Mueller et al., 2024). Therefore, anonymization and removal of copyrighted content is crucial to credibility tools, even when they serve as foundational or backbone models. Paradigms such as unlearning (Chen and Yang, 2023) or LLM editing could be potential research directions to tackle these issues.
@@ 9.4. Practical deployment
Credibility assessment is naturally very challenging when it comes into the practical deployment. Also by proceeding from our own experience, during a deployment in real-world settings, the performance of trained models many times does not achieve the sufficient thresholds. The root of performance decrease between the offline experiments (on the testing data) vs. in the real-world world scenarios can be traced into the out-of-distribution nature of data (from any of relevant aspects, such as topic, format, language, etc.). This imposes a need for more robust solutions, which are, however, difficult to achieve considering already-discussed limited dataset availability and dynamics of online environments.
Furthermore, there is commonly a lack of mutual understanding. Technical practitioners and media professionals often use different terminology for technical concepts and tasks. This can pose challenges on interpretation of AI-predicted outputs, understanding limitations of AI models or interpretation of task results. As a similar challenge, the prior datasets and benchmarks were mostly created by computational linguists or computer scientists, while media professionals were not actively involved in labeling. This may easily result into different perspectives what is credible or not; or where some credibility signals should be detected as present and where not. Additionally, as seen from the survey results, some tasks are already subjective such as bias detection and check-worthy detection tasks. These issues can result in model outputs that can fail to align with the perspectives of media professionals, or only succeed in identifying cherry-picked (unambiguous) examples.
Although it is desirable (from the research perspective) to use open-source models for transparency and reproducibility, hosting of several models can be costly for media organizations or academic institutions. Also researchers, when deploying their models, must consider a dilemma regarding the size of the models. Larger language models can typically achieve a higher classification accuracy. At the same time, they would need to be deployed at costly GPU-machines (in order to make inference feasible and fast enough). As a result, a compromise between computation costs and acceptable performance drop must be found. While methods on distilling large models to smaller version can become helpful in some cases, the performance of distilled models can again be worse in comparison to their original versions.
Last but not least, end users (especially general public), even together with their peers, lack the capability to determine the correctness of AI-based credibility indicators (Lu et al., 2022). This implies serious risks of people being misled by AI-based credibility indicators that are wrong, and consequently believing in or even spreading misinformation. For this reason, already mentioned explainability/interpretability as well as clear communication of expected accuracy of the AI-based systems to end users represent a critical aspect of the whole solution.
@@ 9.5. Multimodal approaches
When tackling credibility assessment, the usage of multimodal approaches grows accordingly to the adjacent increasing complexity of the platforms and the evolution of the communication methods. While text is still one of the most common methods to spread information, it can also be combined with other media types, such as images, videos and audios. Multimodal techniques integrate various forms of data, offering a richer understanding of the content and its credibility signals. The main potential of these methods lies in their ability to overcome limitations inherent in text-only analysis. While Natural Language Processing excels at detecting linguistic patterns that may signal misinformation (e.g., sentiment analysis, linguistic complexity, semantic coherence), it can overlook cues that are often found in non-textual elements. Additionally, as previously commented in Section 8.7, there are promising results of non-NLP methodologies to extract information and perform analysis in the credibility context. Thus, having proven the potential of NLP, non-NLP techniques can be combined with it at the same time.
In this regard, the more the interconnectivity of the media formats and the evolution of online sources (such as newspapers, social networks, blogs and other types of communication frameworks), the more interest in using multimodal approaches exists. Non-credible content often spans beyond text, employing visual, auditive, and interactive elements to manipulate perceptions. Studies have shown that combining textual analysis with multimedia analysis leads to significant improvements in detecting and understanding misinformation (Choudhary et al., 2024; Alam et al., 2022a; Hangloo and Arora, 2022).
Multimodal systems rely on complementary features extracted from text alongside non-NLP methods discussed earlier. For example, low credibility content such a fake news article can back-up its misleading text with an altered image or video to enhance its perceived credibility. In such cases, the integration of visual and textual analysis is critical when flagging the content as suspicious.
Multimodal approaches involve processing multiple data streams simultaneously, combining traditional NLP methods with analysis of visual, auditive, and interactive content to form a wider comprehension of credibility. This integration is typically achieved through ensemble models or architectures that unify the outputs of various sub-models focused on different media types. A common method involves using transformer-based models such as BERT or GPT for text analysis alongside convolutional neural networks (CNNs) for image or video analysis. These models work in tandem, the text models identifying linguistic signals of credibility while the CNNs assess visual features like image authenticity, quality, and relevance. Additional layers of analysis might focus on audio quality, transcription accuracy, or even the relationship between the text and accompanying media (e.g., whether an image truly matches the content described in an article). For example, a multimodal model might analyze a tweet’s text for sentiment and coherence, while also assessing the attached image’s metadata and authenticity. If the image shows signs of manipulation or does not align with the textual claim, the credibility for that piece of content would be lowered.
Even in multimodal systems, textual analysis remains a central component. Text provides rich semantic information, often containing the explicit claims or arguments being made, which serve as the basis for evaluating credibility. Without text, it would be difficult to discern the specific intent behind multimedia content and a more specific tool would be needed to perform the evaluation under certain conditions. Thus, NLP techniques like semantic parsing, sentiment analysis, and entity recognition form the ground of these multimodal approaches, supplying the primary framework for understanding the broader context in which non-textual signals are embedded.
Studies in multimodal credibility assessments consistently show that while non-textual cues (e.g., image authenticity, video manipulation detection) significantly enhance the detection process, they are most effective when paired with deep text analysis. For instance, recent work on fake news propagation graphs combines user engagement patterns with textual signals and visual content to identify disinformation (Qureshi and Malick, 2023). Another promising approach is the integration of social media metadata (such as virality metrics and user behavior) alongside textual content to detect the spread of fake news (Zade et al., 2023; Spezzano et al., 2021). The application of multimodal techniques in credibility assessments is expanding rapidly, particularly in social media, news aggregation platforms, and user-generated content sites like YouTube, Twitter, and Reddit.
@@
In future developments, there is potential to enhance multimodal models by integrating additional modalities, such as biometric signals (e.g. eye-tracking data), interactive content quality, or real-time user engagement metrics. Research on multimodal disinformation detection is ongoing, with several promising methods emerging, such as ensemble techniques that combine the strengths of multiple machine learning models across different modalities (Hansen et al., 2020; Al-alshaqi et al., 2024).

### v2

@@ 9.3. Dataset availability and multilinguality
Dataset availability heavily differs across various categories of credibility signals. Firstly, we observed a significant lack of large-enough datasets providing overall (expertly-determined) annotation of credibility that can be used for training and evaluating solutions on credibility assessment task (Section 4.1). Furthermore, datasets containing annotations for overall credibility and (at least some) credibility signals at the same time are even more scarce. As a result, many researchers opted to use (more available) fake news datasets as a replacement. At this place, we would like to highlight again that fake news annotations cannot reliably replace credibility annotations (non-credible content is not necessarily only false content and vice versa, since credibility is rather a parallel and complementary dimension to veracity; see Section 2).
@@
On the other hand, the situation with dataset availability for individual categories of credibility signals is much better. This is especially thanks to data challenges (particularly SemEval tasks and CLEF CheckThat! Labs as evidenced in Sections 5, 6, or 7) in which either new datasets were introduced or existing datasets were extended (with more data, new languages, or additional types of annotations). Unfortunately, in some cases (e.g., persuasion techniques dataset by Piskorski et al. (2023) introduced in Section 6), the datasets from data challenges are not published completely and a hidden/testing set is not shared with researchers not even when the competition is over (and thus only the training/validation sets remain available for further research).
@@
Especially for international news or the investigation of global claims, journalists and fact-checkers need to verify the credibility of information by cross-checking sources in different languages. Furthermore, emerging disinformation in one country could be spread to other countries, especially when its topic is global (e.g., pandemic, wars, international relations). Therefore, it is important to implement credibility analysis tools that support multiple languages. However, due to the scarcity of multilingual datasets, trained models can exhibit biases towards some languages and cultures and hence can underperform on content in low-resourced languages (as is for example the case for fact-checked claims retrieval as evidenced in (Vykopal et al., 2025b); see Section 7).
In Table 11, we present a summarised overview of language-specific datasets available for different categories of credibility signals. As in many other areas of NLP, English remains the dominant language, with the majority of available resources focused on English-language content. Nevertheless, notable datasets also exist for other widely spoken languages, particularly Arabic and Spanish. In contrast, there is a clear lack of annotated resources for many low- and mid-resource languages, such as Czech or Tamil. Among the covered categories, fact-checked claims demonstrate the broadest language coverage. This can be attributed to the global presence of fact-checking organizations, which produce multilingual artifacts that serve as the foundation for constructing such diverse and multilingual datasets. On the other hand, categories such as factuality and bias show limited availability of resources beyond English, highlighting a significant gap in the development of multilingual tools for credibility signals detection.
| Language | Credibility Signals | |||||||||||||||||||
| Subjectivity | Factuality | Bias |
| English |
(Qian et al., 2022, 2019; Li et al., 2024c) |
| Chinese | – | (Qian et al., 2022, 2019) | – | – | – | (Nielsen and McConville, 2022; Pikuliak et al., 2023) | ||||||||||||||
| German | (Atalla et al., 2011; Piskorski et al., 2023; Banea et al., 2010) | – | (Piskorski et al., 2023; Aksenov et al., 2021) | (Piskorski et al., 2023) | – | (Nielsen and McConville, 2022; Pikuliak et al., 2023) | ||||||||||||||
| Urdu | (Wiebe and Riloff, 2005) | – | – | – | – | (Nielsen and McConville, 2022; Pikuliak et al., 2023) | ||||||||||||||
| Arabic | (Mourad and Darwish, 2013; Banea et al., 2010) | – | – | (Alam et al., 2022b; Hasanain et al., 2023; Almotairy et al., 2024) |
(Nakov et al., 2022b; Nielsen and McConville, 2022; Pikuliak et al., 2023) | ||||||||||||||
| French | (Piskorski et al., 2023; Banea et al., 2010) | – | (Piskorski et al., 2023) | (Piskorski et al., 2023) | – | (Nielsen and McConville, 2022; Pikuliak et al., 2023) | ||||||||||||||
| Polish | (Piskorski et al., 2023) | – | (Piskorski et al., 2023) | (Piskorski et al., 2023) | – | (Nielsen and McConville, 2022; Pikuliak et al., 2023) | ||||||||||||||
| Italian | (Piskorski et al., 2023) | – | (Piskorski et al., 2023) | (Piskorski et al., 2023; Macagno, 2022) | – | (Nielsen and McConville, 2022; Pikuliak et al., 2023) | ||||||||||||||
| Russian | (Piskorski et al., 2023) | – | (Piskorski et al., 2023) | (Piskorski et al., 2023) | – | (Nielsen and McConville, 2022; Pikuliak et al., 2023) | ||||||||||||||
| Spanish | (Banea et al., 2010) | – | – | – |
(Martín et al., 2022; Singh et al., 2024; Nielsen and McConville, 2022; Pikuliak et al., 2023) | ||||||||||||||
| Romanian | (Banea et al., 2010) | – | – | – | – | (Pikuliak et al., 2023) | ||||||||||||||
| Portuguese | (Jeronimo et al., 2020) | – | – | (Macagno, 2022) | – | (Singh et al., 2024; Nielsen and McConville, 2022; Pikuliak et al., 2023) | ||||||||||||||
| Dutch | (Maks and Vossen, 2012) | – | – | – | (Nakov et al., 2022a; Barrón-Cedeño et al., 2024) | (Nielsen and McConville, 2022; Pikuliak et al., 2023) | ||||||||||||||
| Czech | – | – | – | (Baisa et al., 2019) | – | (Nielsen and McConville, 2022; Pikuliak et al., 2023) | ||||||||||||||
| Turkish | – | – | – | – | (Kartal and Kutlu, 2020; Nakov et al., 2021, 2022a) | (Nielsen and McConville, 2022; Pikuliak et al., 2023) | ||||||||||||||
| Hindi | – | – | – | – | (Dhar and Das, 2021) | (Kazemi et al., 2021; Singh et al., 2024; Nielsen and McConville, 2022; Pikuliak et al., 2023) | ||||||||||||||
| Bengali | – | – | – | – | (Dhar and Das, 2021) | (Kazemi et al., 2021; Nielsen and McConville, 2022; Pikuliak et al., 2023) | ||||||||||||||
| Norwegian | – | – | – | – | (Sheikhi et al., 2023) | (Nielsen and McConville, 2022; Pikuliak et al., 2023) | ||||||||||||||
| Bulgarian | – | – | – | – | (Nakov et al., 2021, 2022a) | (Pikuliak et al., 2023) | ||||||||||||||
| Tamil | – | – | – | – | – | (Kazemi et al., 2021; Nielsen and McConville, 2022; Pikuliak et al., 2023) | ||||||||||||||
| Malayalam | – | – | – | – | – | (Kazemi et al., 2021; Nielsen and McConville, 2022; Pikuliak et al., 2023) | ||||||||||||||
| Additional languages | – | – | – | – | – | + 22 (Nielsen and McConville, 2022), + 18 (Pikuliak et al., 2023) | ||||||||||||||
@@
To overcome the scarcity of multilingual datasets, global collaborations could be initiated for creating multilingual resources. In this direction, we can already observe a positive trend within data challenges. Many of them introduced multilingual datasets, commonly considering some languages as surprise ones (i.e., languages that are present in a test set, but missing in a train/validation set). Such approach motivates participants to develop multilingual solutions that are capable to make a prediction in a zero-shot setting (considering a language a predicted content is written in).
@@
Besides datasets availability and multilinguality, a quality of annotations remain another challenge. Annotation of overall credibility as well as individual credibility signals is many times highly subjective (as also shown in (Bhuiyan et al., 2020)), especially in cases such as persuasion techniques where presence of a signal and borders between various signals can be blurred. The situation is getting even more challenging in multilingual settings, where typically native speakers are needed to annotate data. Firstly, acquiring human experts fluent in several languages is challenging itself. Secondly, organizing and consolidating annotation process (including post-annotation verification) is a complex task. Considering also our own experience working with such datasets, the provided labels cannot be easily verified and many times we identified (inevitable) incorrect labels.
@@
Last but not least, credibility assessment and automatic detection of credibility signals naturally happen in very dynamically evolving (online) environment. New topics and global events constantly emerge, causing significant data and concept drifts. In some cases, such drifts can cause that the existing datasets may become obsolete and non-representative. Secondly, the list of credibility signals itself evolves. We can take a machine-generated text as an illustrative example. This kind of signal become highly relevant only recently with generative LLMs becoming easily available for a large end user base (and unfortunately also bad actors). Such new/redefined signals thus naturally result into a demand for new datasets. Finally, the dynamics of this area also lies in its adversarial character. Bad actors (e.g., the ones who are spreading propaganda or disinformation) will always try to get their content undetected as low-credible one by employing various obfuscation techniques. This must be taken into consideration when introducing new datasets. By continuing with an illustrative example of machine-generated text, there are already datasets (e.g., (Macko et al., 2024)) providing besides machine-generated texts also their alternative versions after applying several authorship obfuscation techniques, which allows to train and evaluate more robust detection models.
@@ 9.4. Ethical and legal issues
Credibility assessment of online (primarily social media) content is from its nature an area that must address several ethical and legal issues.
First, such ethical and legal issues are especially prominent when the researched outcomes get deployed and used in practice by end users. The challenge for tool makers is to be fully transparent about the limitations of their tools, to provide guidelines to avoid misleading their users (e.g., by false positives and hallucinations) and to support human control (in line with the human-in-the-loop approach), for more details see Section 9.5.
@@
Second, similarly to other related research areas (e.g., false information detection), credibility assessment methods may be potentially misused by bad actors in order to create content that appears to be more credible. Also by following open-research spirit, publishing credibility assessment systems can theoretically result into misusing such system in the adversarial manner to tune disinformation/propaganda generation systems and allowing them to stay undetected. This kind of potential threat is, however, an analogical issue to the security domain and the principle of security by obscurity. Nothing prevents bad actors to develop their own credibility assessments systems and apply then in adversarial training scenario. Moreover, positive outcomes of credibility assessment research are more tangible, with many practical tools (as also showed across this survey) already put into the hands of media professionals or general public.
@@
Besides potential misuse, additional ethical considerations must be addressed thoroughly during the research activities. First of all, training various classification/detection systems is inherently a subject of potential biases. Such biases can come directly from the datasets used for the training purposes (in terms of data selection, data pre-/post-processing, or data annotation itself; see Section 9.3) or can be introduced during training the models, especially when fine-tuning pre-trained LLMs that have incorporated biases by themselves (including biases between high- and low-resource languages; see Section 9.2). Secondly, the authors should always clearly formulate intended use and failure cases of the trained models/deployed tools (e.g., by means of model cards). In this way, we may prevent media professionals to over-rely on the predicted (potentially incorrect) values. Fortunately, the two-step approach to credibility assessment (i.e., detect more granular credibility signals and then aggregate them) makes the whole process more transparent. Last but not least, explainability and interpretability of the models’ predictions plays a crucial role in this area, since credibility assessment must be credible itself, otherwise it would not provide expected level of trust to its end users.
@@
Besides ethical issues, the research in this area must address also several legal issues. Many of them are shared with other research works on social media. Firstly, as we have also witnessed recently, social media platforms can change their data processing policies and restrict access to their data, which may delay progress in research and development of credibility tools. Additionally, social media data must be anonymized to protect user privacy before being used as training data or for the model inference. Media organizations also impose limitations due to copyright laws, with some not permitting their content to be used in AI tool development. LLMs, especially closed LLMs such as ChatGPT and GPT-4, lack transparency regarding pre-training data and LLMs can memorize content in their pretraining data (Karamolegkou et al., 2023; Mueller et al., 2024). Therefore, anonymization and removal of copyrighted content is crucial to credibility tools, even when they serve as foundational or backbone models. Paradigms such as unlearning (Chen and Yang, 2023) or LLM editing could be potential research directions to tackle these issues.
@@ 9.5. Practical deployment
Deploying credibility assessment and credibility signal detection models in real-world settings presents significant challenges. For media professionals, editorial guidelines in most newsrooms and fact-checking organizations recommend treating AI-based credibility assessment or fact-checking tools as exploratory aids rather than definitive sources of evidence. Fact-checkers are generally advised to use AI-generated outputs as complementary cues – only after accumulating sufficient independent evidence pointing to the falsity of the content in question. A similar critical challenge concerns the general public, who often lacks the expertise to critically assess AI-detected credibility signals (Lu et al., 2022). This poses serious risks – if such indicators are incorrect or poorly communicated, users may be misled, potentially reinforcing belief in or further disseminating false information. These practical concerns stem from multiple underlying (not only purely technological) challenges.
First, explainability, interpretability, and transparent communication of model confidence and limitations are essential components of any credibility assessment system intended for real-world use for both types of end users – media professionals as well as the general public. Many classifiers (e.g., on subjectivity or persuasion techniques) are providing clues that the editor can proofread, maintaining editorial control on the analysis. At the same time, the epistemic shift of the AI-generated content is that humans are now struggling to determine if the content is synthetic or not, and therefore tend to rely more on automated tools, on which exercising editorial control is much more complicated. How can an editor trust an AI-based detector that has an unexplained and non-negligible known rate of false positives? How can she take the reputation risk of writing that the content is synthetic, especially in the context of political life, elections, and public debate, if it proves to be authentic later on? To this end, end users must be able to understand the strengths and weaknesses of models, and model providers must disclose the constraints and limits of the tools they provide (including information on the provenance of the datasets and the training of the models).
The previous research showed that appropriate visual explanations foster end users’ trust in AI-predicted classification labels (Yang et al., 2020). In this direction, Przybyła and Soto (2021) proposed a set of interactive visualizations designed to explain the rationale behind automatic credibility assessments, with the aim of increasing users’ confidence in the underlying methods. In a user study involving 14 participants, the authors found that the stylometric classifier was perceived as more interpretable than the neural classifier, although the latter achieved higher predictive performance. Participants were also significantly more accurate and more confident in their credibility judgments after interacting with the visual explanations. These findings underscore that providing meaningful explanations and interpretability – beyond a simple black-box credibility label or score – remains an open challenge. Moreover, the study highlights that addressing this issue is non-trivial, particularly given that state-of-the-art models such as deep neural networks often suffer from limited inherent explainability, despite their high performance.
Another critical challenge lies in the communication gap between technical practitioners and end users, particularly media professionals and the general public. These groups often operate using distinct terminologies and conceptual frameworks, which can hinder the interpretation of AI-generated outputs or the understanding of AI system limitations. Furthermore, many existing datasets and benchmarks have been developed primarily by computational linguists or computer scientists, without active involvement of end users. This can lead to divergent perspectives regarding what constitutes credible content or when a credibility signal is present. As our survey shows, certain credibility signal detection tasks, such as bias detection (Section 5) and check-worthiness assessment (Section 7), are inherently subjective, which may further exacerbate these discrepancies. Consequently, model predictions may align poorly with end users’ expectations or only succeed on simple unambiguous examples.
Second, detection models making use of credibility signals should minimize false positive rates as much as possible for integration in media organisations and journalism toolkits. However, based on both literature and our own practical experience, we observe that model performance often deteriorates when transitioning from offline evaluations to real-world applications. This discrepancy largely stems from the out-of-distribution (OOD) nature of deployment data, which may differ from training and testing datasets in various aspects, such as topic, format or language. These shifts highlight the need for more robust and generalizable models. However, achieving such robustness is complicated by the already-discussed limited availability of diverse datasets (Section 9.3) and the dynamic nature of online environments.
Furthermore, deployed methods and tools for automatic credibility assessment are inherently exposed to potential adversarial threats. Malicious actors who intentionally disseminate non-credible information may attempt to evade detection by slightly altering the content – creating adversarial examples that exploit vulnerabilities in detection models and trigger incorrect classifications (Przybyła et al., 2024). To be effective in real-world applications, automatic classifiers must therefore be robust to such adversarial manipulations. Recent studies have demonstrated that LLMs, when combined with diversity incentives techniques, can be effectively leveraged for data augmentation (Cegin et al., 2024). By generating diverse paraphrases and content variations, such approaches not only improve model generalization but also increase resilience against potential adversarial attacks – including those generated using similar LLM-based strategies.
From a technical deployment perspective, resource constraints present additional barriers. While from the research perspective, open-source models are desirable for their transparency and reproducibility, hosting and running multiple models can be financially burdensome for media organizations or academic institutions. Researchers and practitioners often face a trade-off between model size and deployment feasibility: larger language models generally yield higher classification performance but require costly GPU infrastructure to support real-time inference. This necessitates compromises between computational cost and acceptable performance degradation. Techniques such as model distillation, which transfer knowledge from large models into smaller more efficient versions, may alleviate some of these challenges, but often at the cost of reduced performance.
@@ 9.6. Multimodal approaches
While text is still one of the most common methods to spread information, in practice, it is commonly combined with other media types, like images, videos and audios. The demand for multimodal approaches also grows with a continuous shift of social media platforms towards short multimedia formats, such as short videos, reels or slideshows. While Natural Language Processing (NLP) allows to detect linguistic patterns that may signal low credibility content (as shown throughout this survey), it can overlook cues that are often found in such non-textual elements. For example, low-credibility content can back-up its misleading text with an altered or fully generated image or video to enhance its perceived credibility.
As previously discussed in Section 8.7, there are promising results on multimodal approaches that involve processing multiple data streams simultaneously, offering a richer understanding of the content and its credibility. The main potential of these methods lies in their ability to overcome limitations inherent in text-only analysis. To this end, they combine traditional NLP methods with analysis of visual, auditive, contextual or interactive content. Even in such multimodal systems, textual analysis still remains an important component as text provides rich semantic information, often containing the explicit claims or arguments being made. Without appropriate text analysis, it would be difficult to discern the specific intent behind multimedia content.
@@
In the future, there is potential to enhance multimodal models by integrating additional modalities, such as biometric signals (e.g., eye-tracking data (Hansen et al., 2020)), interactive content quality, or real-time user engagement metrics. Research on multimodal credibility assessment is therefore still ongoing and provides several promising directions for future research.